PREPARATION OF PAPERS FOR JACoW CONFERENCES^{∗}

J. Poole, C. Petit-Jean-Genaz, CERN, Geneva, Switzerland C. Eyberger^{†}, ANL, Argonne, IL 60439, USA

Abstract

Many conference series have adopted the same standards for electronic publication and have joined the Joint Accelerator Conference Website (JACoW) collaboration [1] for the publication of their proceedings. This document describes the common requirements for the submission of papers to these conferences. Please consult individual conference information for page limits, method of electronic submission, etc. It is not intended that this should be a tutorial in word processing; the aim is to explain the particular requirements for electronic publication at these conference series.

SUBMISSION OF PAPERS

Each author should submit the PostScript and all of the source files (text and figures), to enable the paper to be reconstructed if there are processing difficulties.

MANUSCRIPTS

Templates are provided for recommended software and authors are advised to use them. Please consult the individual conference help pages if questions arise.

General Layout

These instructions are a typical implementation of the requirements. Manuscripts should have:

• Either A4 (21.0 cm × 29.7 cm; 8.27 in × 11.69 in) or US letter size (21 cm × 27.9 cm; 8.5 in × 11.0 in) paper.

• Single spaced text in two columns of 82.5 mm (3.25 in) with 5.0 mm (0.2 in) separation.

• The text located within the margins specified in Table 1 to facilitate electronic processing of the PostScript file.

Table 1: Margin specifications

Margin A4 Paper US Letter Paper
Top  37 mm  19 mm (0.75 in)
Bottom  19 mm  19 mm (0.75 in)
Left  20 mm  20 mm (0.79 in)
Right  20 mm  26 mm (1.02 in)

The layout of the text on the page is illustrated in Fig. 1. Note that the paper’s title and the author list should be the width of the full page.Tables and figures may span the whole 170 mm page width, if desired (see Fig. 2), but fullwidth figures should be placed at either the top or bottom of a page to ensure a proper flow of the text (Word only).

Figure 1: Layout of papers.

Fonts

In order to produce good Adobe Acrobat PDF files, authors using a L^{A}TEX template are asked to use only Times (in roman (standard), bold or italic) and symbols from the standard PostScript set of fonts. In Word use only Symbol and, depending on your platform, Times or Times New Roman fonts in standard, bold or italic form.

Title and Author List

The title should use 14 pt bold uppercase letters and be centred on the page. Individual letters may be lowercase to avoid misinterpretation (e.g., mW, MW). To include a funding support statement, put an asterisk after the title and a footnote at the bottom of the first column on page 1; in L^{A}TEX use \thanks.

The names of authors, their organisations/affiliations and mailing addresses should be grouped by affiliation and listed in 12 pt upper and lowercase letters. The name of the submitting or primary author should be first, followed by the co-authors, alphabetically by affiliation.

∗ Work supported by ...

† cee@aps.anl.gov

Figure 2: Example of a full-width figure showing the JACoW Team at their annual meeting in 2008. The figure carries a multi-line caption which has to be justified, rather than centred.

Section Headings

Section headings should not be numbered. They should use 12 pt bold uppercase letters and be centred in the column. All section headings should appear directly above the text – there should never be a column break between a heading and the following paragraph.

Subsection Headings

Subsection headings should not be numbered. They should use 12 pt italic letters and be left aligned in the column. Subsection headings should appear directly above the text – there should never be a column break between a heading and the following paragraph.

Paragraph Text

Paragraphs should use 10 pt font and be justified (touch each side) in the column. The beginning of each paragraph should be indented approximately 3 mm (0.13 in). The last line of a paragraph should not be printed by itself at the beginning of a column nor should the first line of a paragraph be printed by itself at the end of a column.

Figures, Tables and Equations

Place figures and tables as close to the place of their mention as possible. Lettering in figures and tables should be large enough to reproduce clearly. Use of non-approved fonts in figures often leads to problems when the files are processed. L^{A}TEX users – please be sure to use non bitmapped versions of Computer Modern fonts in equations (type 1 PostScript fonts are required and their use is described in the JACoW help pages [2]).

All figures and tables must be given sequential numbers (1, 2, 3, etc.) and have a caption (10 pt font) placed below the figure or above the table being described. Captions that are one line should be centred in the column, while captions that span more than one line should be justified. The L^{A}TEX template uses the ’booktabs’ package to format the tables. A simple way to introduce figures into a Word document is to place them inside a table which has no borders. This is done in Word as follows:

• Insert a continuous section break.

• Insert two empty lines (will make subsequent editing easier).

• Insert another continuous section break.

• Click between the two section breaks and Format → columns → Single.

• Table → Insert single column, two row table.

• Paste the figure in the first row and adjust the size as appropriate.

• Paste/Type the caption in the second row and apply figure caption style.

• Table → Table properties → Borders and shading → None.

• Table → Table properties → Alignment → Center.

• Table → Table properties → Text wrapping → None.

• Remove the blank lines from in and around the table.

• If necessary play with the cell spacing and other parameters to improve appearance.

If a displayed equation needs a number, place it flush with the right margin of the column (see Eq. 1). The equation itseld should be centred, if possible. Units should be written using the roman (standard) font, not the italic font.

C_{B} = \frac{q^{3}}{3ϵ_{0}mc} = 3.54 μeV/T  (1)

References

All bibliographical and web references should be numbered and listed at the end of the paper in a section called “References”. When referring to a reference in the text, place the corresponding reference number in square brackets [3]. A URL may be added as part of a reference, but its hyperlink should NOT be added.

Footnotes

Footnotes on the title and author lines may be used for acknowledgements, affiliations and e-mail addresses. A nonnumeric sequence of characters (*, †, ‡, §) should be used. All other notes should be included in the references section and use the normal numeric sequencing.

Word users–do not use Word’s footnote feature (Insert, Footnote) to insert footnotes as this will create formatting problems. Instead, insert footnotes manually in a text box at the bottom of the first column with a line at the top of the text box to separate the footnotes from the rest of the paper’s text. The easiest way to do this is to copy the text box from the JACoW template and paste it into your own document. These ’pseudo footnotes’ in the text box should only appear at the bottom of the first column on the first page.

Acronyms

Acronyms should be defined the first time they appear.

STYLES

Table 2 summarises the fonts and spacings used in the styles of a JACoW template (these are implemented in the L^{A}TEX class file).

PAGE NUMBERS

DO NOT have any page numbers. They will be added when the final proceedings are produced.

TEMPLATES

Templates and examples can be retrieved through Web browsers like Firefox and Internet Explorer by saving to disk. See your local documentation for details of how to do this.

Template documents for the recommended word processing software are available from the JACoW Website and exist for L^{A}TEX, Microsoft Word (Mac and PC) and OpenOffice for US letter and A4 paper sizes.

Use the correct templates for your paper size and version of Word. Do not transport Microsoft Word documents across platforms, e.g. Mac ↔ PC.

Table 2: Summary of Styles

Style Font Space Space
Before After
Title  14 pt  0 pt 3 pt

Upper case except for required lower case letters Bold

Author list 12 pt  9 pt 12 pt

Upper and Lower case

Section  12 pt  9 pt 3 pt
heading  Uppercase

bold

Subsection 12 pt  6 pt 3 pt
heading  Initial Caps

Italic

Figure  10 pt  3 pt 6 pt

Captions

Table  10 pt  3 pt 3 pt

Captions

Equations 10 pt base font  12 pt 12 pt
References 10 pt, justified with  0 pt 0 pt

7 mm hanging indent

Please see the help files for instructions on how to install templates in your Microsoft templates folder.

CHECKLIST FOR ELECTRONIC PUBLICATION

• Use only Times or Times New Roman (standard, bold or italic) and Symbol fonts for text – 10 pt minimum except References which can be 9 pt or 10 pt.

• Figures should use Times or Times New Roman (standard, bold or italic) and Symbol fonts when possible – 6 pt minimum.

• Check that the PostScript file prints correctly.

• Check that there are no page numbers.

• Check that the margins on the printed version are within ±1 mm of the specification.

• L^{A}TEX users can check their margins by invoking the boxit option.

REFERENCES

[1] C. Petit-Jean-Genaz and J. Poole, “JACoW, A service to the Accelerator Community”, EPAC’04, Lucerne, July 2004,

                                                        THZCH03, p. 249, http://www.JACoW.org.

[2] A. Name and D. Person, Phys. Rev. Lett. 25 (1997) 56.

[3] A.N. Other, “A Very Interesting Paper”, EPAC’96, Sitges, June 1996, MOPCH31, p. 7984 (1996),

                                                        http://www.JACoW.org.